# The signless Laplacian spectral Turán problems for hypergraphsthanks: Research was partially supported by the National Nature Science Foundation of China (grant numbers 12331012, 12571375)

### v1

@@ Abstract
Let $\mathcal{H}=(V,E)$ be an $r$-uniform hypergraph on $n$ vertices. The signless Laplacian spectral radius of $\mathcal{H}$ is defined as the maximum modulus of the eigenvalues of the tensor $\mathcal{Q}(\mathcal{H})=\mathcal{D}(\mathcal{H})+\mathcal{A}(\mathcal{H})$, where $\mathcal{D}(\mathcal{H})$ and $\mathcal{A}(\mathcal{H})$ are the diagonal tensor of degrees and adjacency tensor of $G$, respectively. In this paper, we establish a general theorem that extends the spectral Turán result of Keevash, Lenz and Mubayi [SIAM J. Discrete Math., 28 (4) (2014)] to the setting of signless Laplacian spectral Turán problems. We prove that for any family $\mathcal{F}$ of $r$-uniform hypergraphs that is degree-stable with respect to a family $\mathcal{H}_{n}$ of $r$-uniform hypergraphs and whose extremal constructions satisfy certain natural assumptions, the signless Laplacian spectral Turán problem can be effectively reduced to the corresponding problem restricted to the family $\mathcal{H}_{n}$. As a concrete application, we completely characterize the signless Laplacian spectral extremal hypergraph for the Fano plane.
@@ 1 Introduction
Let $\mathcal{A}$ be an order $r\geqslant 2$, dimension $n$ tensor, $\bm{x}=(x_{1},x_{2},\ldots,x_{n})^{\mathrm{T}}$ be a column vector of dimension $n$. Then $\mathcal{A}\bm{x}$ is a vector in $\mathbb{C}^{n}$, whose $i$-th component is as the following
@@
We now list some classical results on spectral Turán problems of graphs. In 1986, Wilf [33] demonstrated that if $G$ is an $n$-vertex $K_{k+1}$-free graph, then $\lambda(G)\leq(1-1/k)n$. Motivated by this, Nikiforov [23] and Guiduli [11] independently proved that such a graph $G$ must satisfy $\lambda(G)\leq\lambda\left(T_{n,k}\right)$, thereby extending Turán’s Theorem to the spectral version. In 2013, Abreu and Nikiforov [4] proved that if $G$ is an $n$-vertex $K_{k+1}$-free graph, then $q(G)\leq\left(1-1/k\right)2n$. Furthermore, He, Jin, and Zhang [12] showed that such a graph $G$ satisfies $q(G)\leq q\left(T_{n,k}\right)$, thereby obtaining a spectral version of Turán’s theorem in terms of the signless Laplacian spectral radius. Let $F$ be a color-critical graph with $\chi(F)=r+1$. For sufficiently large $n$, Simonovits [30] determined the Turán number of $F$, proving that the unique extremal graph on $n$ vertices is $T_{n,r}$. Nikiforov [24] established a spectral counterpart to this result, showing that $T_{n,r}$ attains the maximum adjacency spectral radius among all $F$-free graphs on $n$ vertices. Recently, Zheng, Li, and Li [36] settled the signless Laplacian spectral extremal problem for $F$, determining the maximum signless Laplacian spectral radius explicitly. For more related information, we refer the reader to [12, 17, 25, 32, 35].
@@
Despite these advances, our understanding of the signless Laplacian spectral Turán problem of hypergraphs remains relatively limited. In this paper, we establish a criterion for the signless Laplacian spectral Turán problem of hypergraphs, as stated in Theorem 2.1. Furthermore, as an application of Theorem 2.1, we determine the maximum signless Laplacian spectral radius among all Fano plane-free $3$-graphs on $n$ vertices.
@@ 3 Preliminaries
Moreover, equality $q(\mathcal{G})=q(\mathcal{H}_{n})$ holds only if $\mathcal{G}\in\mathcal{H}_{n}$. Furthermore, since $\mathcal{B}_{n}$ is the unique maximizer in $\mathfrak{B}_{n}$, and given the inclusion $\mathcal{B}_{n}\in\mathcal{H}_{n}\subseteq\mathfrak{B}_{n}$, it follows that $\mathcal{B}_{n}$ is also the unique maximizer in $\mathcal{H}_{n}$. We thus conclude that $q(\mathcal{G})\leq q(\mathcal{B}_{n})$, with equality if and only if $\mathcal{G}=\mathcal{B}_{n}$. The proof is complete.∎
@@ 6 Concluding remarks
In this paper, we establish a general criterion and apply it to the signless Laplacian spectral Turán problem of Fano plane. Our findings indicate that any hypergraph Turán problem which has the degree stability property and whose extremal construction satisfies some assumptions admits a corresponding signless Laplacian spectral result. In fact, one could push our results further to hypergraphs, such as the expension of complete graphs, the generalized fans, and the cancellative hypergraphs.

### v2

@@ Abstract
Let $\mathcal{H}=(V,E)$ be an $r$-uniform hypergraph on $n$ vertices. The signless Laplacian spectral radius of $\mathcal{H}$ is defined as the maximum modulus of the eigenvalues of the tensor $\mathcal{Q}(\mathcal{H})=\mathcal{D}(\mathcal{H})+\mathcal{A}(\mathcal{H})$, where $\mathcal{D}(\mathcal{H})$ and $\mathcal{A}(\mathcal{H})$ are the degree diagonal tensor and the adjacency tensor of $\mathcal{H}$, respectively. In this paper, we establish a general theorem that extends the spectral Turán result of Keevash, Lenz and Mubayi [SIAM J. Discrete Math., 28 (4) (2014)] to the setting of signless Laplacian spectral Turán problems. We prove that if a family $\mathcal{F}$ of $r$-uniform hypergraphs is degree-stable with respect to a family $\mathcal{H}_{n}$ of $r$-uniform hypergraphs and its extremal constructions satisfy certain natural assumptions, then the signless Laplacian spectral Turán problem for $\mathcal{F}$ can be effectively reduced to the corresponding problem restricted to the family $\mathcal{H}_{n}$. As a concrete application, we completely determine the extremal hypergraph that maximizes the signless Laplacian spectral radius among all Fano plane-free $3$-uniform hypergraphs, showing that the unique extremal hypergraph is the balanced complete bipartite $3$-uniform hypergraph.
@@ 1 Introduction
Let $\mathcal{A}$ be an order $r\geq 2$, dimension $n$ tensor, $\bm{x}=(x_{1},x_{2},\ldots,x_{n})^{\mathrm{T}}$ be a column vector of dimension $n$. Then $\mathcal{A}\bm{x}$ is a vector in $\mathbb{C}^{n}$, whose $i$-th component is as the following
@@
We now list some classical results on spectral Turán problems of graphs. In 1986, Wilf [33] demonstrated that if $G$ is an $n$-vertex $K_{k+1}$-free graph, then $\lambda(G)\leq(1-1/k)n$. Motivated by this, Nikiforov [23] and Guiduli [11] independently proved that such a graph $G$ must satisfy $\lambda(G)\leq\lambda\left(T_{n,k}\right)$, thereby extending Turán’s Theorem to the spectral version. In 2013, Abreu and Nikiforov [4] proved that if $G$ is an $n$-vertex $K_{k+1}$-free graph, then $q(G)\leq\left(1-1/k\right)2n$. Furthermore, He, Jin, and Zhang [12] showed that such a graph $G$ satisfies $q(G)\leq q\left(T_{n,k}\right)$, thereby obtaining a spectral version of Turán’s theorem in terms of the signless Laplacian spectral radius. We say that a graph $F$ is color-critical if there exists an edge $e$ of $F$ such that $\chi(F-e)<\chi(F)$, where FeF-e denotes the graph obtained from $F$ by deleting the edge $e$. Let $F$ be a color-critical graph with $\chi(F)=r+1$. For sufficiently large $n$, Simonovits [30] determined the Turán number of $F$, proving that the unique extremal graph on $n$ vertices is $T_{n,r}$. Nikiforov [24] established a spectral counterpart to this result, showing that $T_{n,r}$ attains the maximum adjacency spectral radius among all $F$-free graphs on $n$ vertices. Recently, Zheng, Li, and Li [36] settled the signless Laplacian spectral extremal problem for $F$, determining the maximum signless Laplacian spectral radius explicitly. For more related information, we refer the reader to [12, 17, 25, 32, 35].
@@
Despite these advances, the signless Laplacian spectral version has received much less attention, partly because of the additional structural information encoded in the degree tensor. In this paper, we establish a criterion for the signless Laplacian spectral Turán problem of hypergraphs, as stated in Theorem 2.1. Furthermore, as an application of Theorem 2.1, we determine the maximum signless Laplacian spectral radius among all Fano plane-free $3$-graphs on $n$ vertices.
@@ 3 Preliminaries
Moreover, equality $q(\mathcal{G})=q(\mathcal{H}_{n})$ holds only if $\mathcal{G}\in\mathcal{H}_{n}$. Furthermore, since $\mathcal{B}_{n}$ is the unique maximizer in $\mathfrak{B}_{n}$, and given the inclusion $\mathcal{B}_{n}\in\mathcal{H}_{n}\subseteq\mathfrak{B}_{n}$, it follows that that $\mathcal{B}_{n}$ is also the unique maximizer in $\mathcal{H}_{n}$. We thus conclude that $q(\mathcal{G})\leq q(\mathcal{B}_{n})$, with equality if and only if $\mathcal{G}=\mathcal{B}_{n}$. The proof is complete.∎
@@ 6 Concluding remarks
In this paper, we establish a general criterion and apply it to the signless Laplacian spectral Turán problem for the Fano plane. Our findings indicate that any hypergraph Turán problem that has the degree stability property and whose extremal constructions satisfy certain assumptions admits a corresponding signless Laplacian spectral result. Potential future research directions include applying this framework to other forbidden configurations such as expansions of complete graphs, generalized fans, and cancellative hypergraphs.